OpenStudy (anonymous):

write an expression to find the ninth term of this sequence then write the next three terms of the sequence. 3,10,17,24,...;n=25

OpenStudy (saifoo.khan):

It's an arithmetic progression.

Parth (parthkohli):

This is an arithmetic sequence with a common difference of 7.

OpenStudy (anonymous):

f(n+1)=f(n)+7

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow a_{n} = a_{1} + (n - 1)d }\)

OpenStudy (anonymous):

the ninth term is 25 and the fourth is 24

OpenStudy (saifoo.khan):

Q:How i came to know? A: as we can see the sequence is constantly increasing with 7 units. so.. \[\Large T_9 = 3 + (9-1)7\]

OpenStudy (anonymous):

wtf i have no clue what you are doing

Parth (parthkohli):

An arithmetic sequence is when you keep adding a same number.

OpenStudy (saifoo.khan):

wtf = wow, that's fantasic.

Parth (parthkohli):

To find the nth term, you have to use this formula. \(\Large \color{MidnightBlue}{\Rightarrow a_{n} = a_{1} + (n - 1)d }\)

OpenStudy (saifoo.khan):

fantastic*

Parth (parthkohli):

In my formula, n is the nth term and d is the common difference.

OpenStudy (anonymous):

tell me a equation i can use please

Parth (parthkohli):

Which I've already told.

OpenStudy (anonymous):

give me the equation with numbers and stuff please

OpenStudy (saifoo.khan):

Patience..

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 3 + 7(n - 1) }\)

Parth (parthkohli):

For example, for the 4th term, just replace n with 4.